OpenStudy (anonymous):

Which function below is the inverse of f(x) = x2 − 16?

OpenStudy (johnweldon1993):

Inverse Means you switch the 'x' and the 'y' in the equation...and then solve for 'y' again... So here y = x^2 - 16 becomes x = y^2 - 16 Now how would you solve for 'y' ?

OpenStudy (anonymous):

Yup pretty self explanatory actually.